# Learning Interior Point Method for AC and DC Optimal Power Flow

### v1

@@ V-C GI-LSTM Performance
The evaluation metrics of GI-LSTM demonstrate its effectiveness. For instance, the loss function of the trained GI-LSTM on two test systems is shown in Fig. 6. We observed that both the training and validation losses converged toward small values, indicating that the GI-LSTM model was effectively optimized and capable of learning meaningful patterns from the data. A high coefficient of determination $R^{2}$ across all four test systems further confirms the model’s predictive accuracy, as it demonstrates a strong alignment between the predicted and actual values in the validation phase. These high-accuracy projections are refined with IPM to improve numerical stability. The performance metrics for each system are summarized in Table III, highlighting the consistency and generalizability of the proposed approach across different system sizes.
@@ V-D GI-LSTM Outpout
The GI-LSTM model predicts outputs that are both accurate and close to the exact solution. For instance, as Fig. 7, the results generated by GI-LSTM closely match those obtained using IPM, which demonstrates the model’s strong predictive capability.
@@
However, in power system operations, ensuring feasibility is critical. Even minor violations of operational limits—such as voltage, power flow, or generation constraints—can pose serious risks to system stability and reliability. To address this, we refine the GI-LSTM predictions by implementing IPM using the GI-LSTM projected central path. This step ensures that the final solution remains almost the same as the optimal OPF solution and strictly satisfies all system physical and operational constraints. Hence, the feasibility of the L-IPM solutions for OPF is guaranteed.
@@ V-E L-IPM
Simulation Results Summary: The key observation from the L-IPM-based OPF evaluation is highlighted below. Table V summarizes the number of iterations required to solve the OPF problem across various load scenarios. We report the minimum and maximum number of iterations observed during the simulations for each test system. As the system size increases, IPM exhibits higher sensitivity to load variations, resulting in a larger gap between the minimum and maximum number of iterations. This behavior is primarily due to more active inequality constraints under stressed conditions. In contrast, such variation is not observed in the number of L-IPM iterations, highlighting its robustness to load fluctuations. The percentage of average iteration reduction in L-IPM compared to classic IPM is shown in Fig 8(c). For instance, in the case of ACOPF on the 2869-bus system, L-IPM reduced the number of iterations by an average of 85.5%, which is a substantial improvement compared to the classical IPM state.
TABLE V: OPF Iteration Comparison using IPM and L-IPM
DCOPF ACOPF Enhancement (%)
min max min max min max
3-bus 10 (4) 10 (4) 11 (4) 12 (4) 60 66.7
24-bus 8 (4) 37 (5) 13 (4) 16 (4) 50 86.5
118-bus 9 (4) 11 (4) 14 (4) 15 (4) 55.6 73.4
2869-bus 11 (4) 34 (5) 25 (7) 55 (8) 63.7 85.5
Values in parentheses indicate the number of L-IPM iterations.
@@
For a more detailed comparison, we evaluated the classical IPM and the proposed L-IPM feasibility and complementarity conditions. As Fig. 10, L-IPM significantly reduces the error to below 10e610e-6 after just one IPM iteration. We analyzed the structure of the Newton steps, particularly the $A$ matrix in the linear system $Ax=b$, and observed that most of the significant changes occur during the early iterations. The remaining iterations primarily refine the solution by accurately identifying violated constraints and adjusting the slack variables within the IPM framework. The L-IPM learns the behavior of these variables, and with a carefully selected set of load scenarios, as described in the data generation section, it develops an understanding of the feasible region of the system, including generation limits and line capacities.
@@
In extreme load cases, we found that L-IPM needs up to two IPM iterations to converge. This suggests that while GI-LSTM outputs may need slightly more effort to handle highly stressed scenarios, it still provides a significantly faster solution compared to solving the problem entirely with the traditional IPM. Even in these challenging cases, convergence was achieved with only a few iterations.
@@
Given the large scale of the system, load scenarios have a significant impact on the final solution. In particular, IPM struggled to find a feasible solution under poorly conditioned load scenarios, even after many iterations. The solver sometimes reached as many as 53 iterations without convergence. As illustrated in Fig. 17(a), we simulated a challenging load scenario where neither IPM nor L-IPM was able to converge. Even in scenarios where no feasible solution exists, L-IPM can reach this outcome more quickly than the standard IPM. When such a poorly conditioned load scenario is input into our proposed method, the grid-informed structure of GI-LSTM helps ensure that its output remains within or close to the feasible region. After the IPM refinement step, we observed that the IPM fails after approximately five iterations, returning a numerical failure error, similar to the behavior seen in the standalone IPM run. The trend of the feasibility error under this scenario is presented in Fig. 17(b), further highlighting how the hybrid approach accelerates the detection of infeasibility while still adhering to system constraints as closely as possible.
@@ VI Conclusion
This paper proposes a feasibility-guaranteed learning interior point method (L-IPM) aiming to accelerate the runtime of AC and DCOPF problems. While early IPM iterations carry valuable information about the optimal solution, the later ones mainly serve to ensure feasibility. Leveraging this observation, we propose an L-IPM that reduces the number of costly iterations. Our method uses a GI-LSTM to project the IPM central path and find a feasible point near the optimal solution. A final IPM refinement stage confirms feasibility and improves solution accuracy.
To evaluate the proposed approach, we tested it on four systems ranging from small (3- and 24-bus) to medium and large (118- and 2869-bus) networks. These case studies demonstrate the scalability, accuracy, and computational efficiency of our method. Results show that L-IPM, when trained on carefully selected load scenarios, significantly accelerates convergence, both in runtime and in the number of IPM iterations. Under challenging load conditions, L-IPM consistently outperformed traditional IPM by requiring fewer iterations. In many cases, it reduced runtime by about 90% compared to the classic IPM. Further analysis of feasibility and complementarity conditions confirmed the robustness of the proposed method, as it rapidly reduced constraint violations while maintaining solution accuracy. Although some extreme load scenarios—particularly in the 2869-bus system—introduced challenges, L-IPM still identified infeasibility more quickly than conventional IPM.

### v2

@@ V-C GI-LSTM Performance
The evaluation metrics of GI-LSTM demonstrate its effectiveness. For instance, the loss function of the trained GI-LSTM on two test systems is shown in Fig. 6. We observed that both the training and validation losses converged toward small values, indicating that the GI-LSTM model was effectively optimized and capable of learning meaningful patterns from the data. In this model, a slightly lower validation loss is expected compared to the training loss. This primarily results from dropout, which is active during training but disabled during validation, making training inherently noisier because random units are temporarily zeroed at each iteration. Consequently, training is performed under a deliberately perturbed network, whereas validation uses the full model capacity. In addition, training loss is averaged over mini-batches during weight updates, whereas validation loss is computed after each epoch using the updated parameters. Overall, this behavior is consistent with stable regularized training and improved generalization.
A high coefficient of determination $R^{2}$ across all four test systems further confirms the model’s predictive accuracy, as it demonstrates a strong alignment between the predicted and actual values in the validation phase. These high-accuracy projections are verified with IPM. The performance metrics for each system are summarized in Table III, highlighting the consistency and generalizability of the proposed approach across different system sizes.
@@ V-D GI-LSTM Outpout
The GI-LSTM model predicts outputs that are both accurate and close to the exact solution. For instance, as Fig. 7, the results generated by GI-LSTM closely match those obtained using IPM, which demonstrates the model’s strong predictive capability. Even minor violations of operational limits—such as voltage, power flow, or generation constraints—can pose serious risks to system stability and reliability. To address this, we sanity check the GI-LSTM predictions by implementing IPM using the GI-LSTM projected central path. This step ensures that the final solution remains the same as the optimal OPF solution and strictly satisfies all system physical and operational constraints. Hence, the proposed GI-LSTM promotes high-quality L-IPM OPF solutions by steering the trajectory toward the feasible region, while a final IPM step verifies solution accuracy.
@@ V-E L-IPM
Simulation Results Summary: Table V reports the minimum and maximum number of iterations required to solve the OPF problem across all load scenarios for each test system. As system size increases, the classical IPM becomes more sensitive to load variations, leading to a larger gap between the minimum and maximum iteration counts, mainly due to more active inequality constraints under stressed conditions. In contrast, the number of L-IPM iterations shows no such variation, highlighting its robustness to load fluctuations. The percentage reduction in average iterations of L-IPM relative to classical IPM is shown in Fig. 8(c); for example, in ACOPF on the 2869-bus system, L-IPM reduces the average number of iterations by 85.5%, representing a substantial improvement.
TABLE V: OPF Speedup Comparison of L-IPM over IPM
Avg. Speedup ($\times$) Enhancement (%)
DC AC min max
3-bus 2.50 2.88 60 66.7
24-bus 4.70 3.63 50 86.5
118-bus 2.50 3.63 55.6 73.4
2869-bus 4.78 5.22 63.7 85.5
Avg. speedup ($\times$) is computed as the average speedup over the analyzed OPF cases for each OPF type; values >1>1 indicate that L-IPM is faster.
@@
For a more detailed comparison, we evaluated the classical IPM and the proposed L-IPM feasibility and complementarity conditions. As Fig. 10, L-IPM significantly reduces the error to below 10e610e-6. We analyzed the structure of the Newton steps, particularly the $A$ matrix in the linear system $Ax=b$, and observed that most of the significant changes occur during the early iterations. The remaining iterations primarily refine the solution by accurately identifying violated constraints and adjusting the slack variables within the IPM framework. The L-IPM learns the behavior of these variables, and with a carefully selected set of load scenarios, as described in the data generation section, it develops an understanding of the feasible region of the system, including generation limits and line capacities.
@@
Infeasible Cases Analysis: Given the large scale of the system, load scenarios have a significant impact on the final solution. In particular, IPM struggled to find a feasible solution under poorly conditioned load scenarios, even after many iterations. The IPM solver sometimes reached as many as 53 iterations without convergence. As illustrated in Fig. 17(a), we simulated a challenging load scenario where neither IPM nor L-IPM was able to converge. Even in scenarios where no feasible solution exists, L-IPM can reach this outcome more quickly than the standard IPM. When such a poorly conditioned load scenario is provided to the proposed method, the grid-informed design of GI-LSTM helps keep the predicted output within, or close to, the feasible region; however, the method correctly flags the solution as infeasible. To confirm this assessment, we examined the subsequent IPM verification step and observed that IPM fails to converge after approximately five iterations, returning a numerical failure error—consistent with the behavior of a standalone IPM run under the same scenario. The trend of the feasibility error under this scenario is presented in Fig. 17(b), further highlighting how the approach accelerates the detection of infeasibility while still adhering to system constraints as closely as possible.
@@
V-F Performance of Direct Solution Prediction vs. Path-Projection
A broad range of machine learning approaches has been explored to approximate OPF solutions by directly predicting dispatch and voltage variables through supervised learning, grid-informed neural networks, reinforcement learning, or graph-based models. In contrast, the proposed method targets the internal behavior of the OPF solver. Specifically, we leverage IPM- a widely used benchmark solver- and learn to project its central-path trajectory using informative early iterations, rather than replacing the optimizer with a black-box predictor.
A natural question is whether one can bypass the iterative process entirely by directly predicting the final OPF solution (or providing an aggressively accurate warm start). To examine the practical value of such direct prediction in the context of IPM, we conduct two stress tests: (i) initializing IPM with the exact optimal solution or 0% neural network (NN), and (ii) initializing IPM with a perturbed solution obtained by applying a $2\%$ uniform random multiplicative error (2% NN).
The corresponding results are summarized in Table VII. It shows that, in some cases, IPM requires up to 50% more iterations to satisfy its stopping criteria, even when initialized at the optimal primal point. This behavior is expected in a primal–dual IPM: convergence is governed not only by the primal variables, but also by the associated slack and dual variables and their complementarity/centrality conditions. Consequently, a highly accurate primal warm start alone does not necessarily place the method near the central path, and may not yield consistent computational savings.
By contrast, the proposed L-IPM framework exploits the early IPM iterations, which already encode rich information about the grid physics, active constraints, and the direction of the central path toward optimality. These early-iteration features enable effective trajectory projection and substantially improve solution quality. Finally, a concluding IPM verification step is used to assess accuracy and practicality for operational use. As shown in Table VII, the proposed method reduces the IPM solve to only 4 iterations across all test cases, leading to a lower runtime than the MATPOWER and NN warm-start baselines.
TABLE VII: Comparison of IPM iteration counts under different warm-start strategies vs. GI-LSTM.
System DCOPF ACOPF
MATPOWER 0% NN 2% NN GI-LSTM MATPOWER 0% NN 2% NN GI-LSTM
3-bus 10 13 13 4 11 16 16 4
24-bus 25 29 30 4 13 17 17 4
118-bus 11 14 15 4 14 18 19 4
2869-bus 12 18 18 4 36 45 46 4
@@ VI Conclusion
This paper proposes a learning interior point method aiming to accelerate the runtime of AC and DC OPF problems. IPM finds the optimal solution by following a central path iteratively, similar to a time series. While early IPM iterations carry valuable information about the optimal solution, later ones mainly serve to ensure feasibility and are typically computationally expensive. Leveraging this, we propose a grid-informed L-IPM that reduces the number of costly iterations required to project the IPM central path and identify the optimal solution, and then verifies its feasibility using a final IPM step.
To evaluate the proposed approach, we tested it on four systems ranging from small (3- and 24-bus) to medium and large (118- and 2869-bus) networks. These case studies demonstrate the scalability, accuracy, and computational efficiency of our method. Results show that L-IPM, when trained on carefully selected load scenarios, significantly accelerates convergence, both in runtime and in the number of IPM iterations. Under challenging load conditions, L-IPM consistently outperformed traditional IPM by requiring fewer iterations. In many cases, it reduced runtime by about 90% compared to the classic IPM.
Further analysis of feasibility and complementarity conditions confirmed the robustness of the proposed method, as it rapidly reduced constraint violations while maintaining solution accuracy. Although some extreme load scenarios—particularly in the 2869-bus system—introduced challenges, L-IPM still identified infeasibility more quickly than conventional IPM. A comparative analysis against classical OPF solution-prediction methods shows that initializing IPM with a high-quality primal prediction does not necessarily reduce the IPM iteration count or runtime, since IPM may still require additional iterations to properly adjust the dual variables (and barrier-related terms) and satisfy its stopping criteria. In contrast, the proposed method leverages the OPF solver’s central-path behavior to provide a more informative direction, thereby reducing the total number of IPM iterations.